A steel rod increases its length by 5mm when the temperature increases by 10°c. What is the initial length of the rod if the coefficient of linear expansion for steel is 0.00011°c
1
Expert's answer
2020-01-28T13:22:51-0500

Here because the rod is heated, the temperature increases the length of the rod and this is known as linear expansion of materials.


From the thermal expansion equation,

"\u0394L = \u03b1 * L * \u0394T"


Change in Length (ΔL)

original length of the rod (L)

change in temperature ( ΔT)

And α is coefficient of linear expansion which depends on the material used.



Substituting,

"5*10^{-3}=0.00011*L*10\\
L=4.54545455m"